In 1960, the average cost of a standard automobile was $1,100. Currently, the cost of a standard automobile in 2020 is $33,500. What is the average cost of an automobile a person can expect to pay in the year 2030 if the cost is assumed to increase in a linear fashion?  
Answer #1

and Cx2, Y2) = ( 2020, 3350o) Here the standard cost in 1960 is $ 1100 2020 it is $ 33,500 since it is increasing in linear F=) 32400 1100 - y 1960-х 32400 C 1960-x) 60 C1100 - y) 58001 2) 540 (1960-x) = c 1100 - 9 1058400 - 54026 = 1100 - y =) 5402C